OpenStudy (anonymous):

Consider the line L defined by y= 1/2x + 1 the line that is perpendicular to L and passes through the point ( 3,0) can be written in slope intercept form as y=? x+ ?

OpenStudy (ybarrap):

slope of perpendicular line = -2 for perpendicular line to given (reusing variable y as you did): y = -2x + b at x =3, y = 0 0 = -2(3) + b b = 6 so equation of line perpendicular to given is y = -2x + 6 Ans. y = -2x + b

OpenStudy (anonymous):

Thanks appreciate your help!!!

OpenStudy (ybarrap):

np